OpenStudy (anonymous):

Solve the equation for x: (4x - 6)(2x - 5) = 0 A. {1.5, 2.5} B. {0, 2.5} C. {5, 6} D. no solution I think he answer is B but i am not for sure

OpenStudy (anonymous):

A

OpenStudy (anonymous):

A is correct, x = 6/4 and x = 5/2 so x = 1.5 and 2.5

OpenStudy (anonymous):

thank you